OpenStudy (anonymous):

PLEASE HELP? What is the value of the y variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −6 answer choices: 6 −3 3 −6

OpenStudy (amistre64):

one way would be to solve each equation for x; and see if the y values give the same x solutions

OpenStudy (anonymous):

I dont get what you mean @amistre64

OpenStudy (amistre64):

the others methods like substitution and elimination work also ..

OpenStudy (amistre64):

if the y values you have for answer choices is true; it will produce the same x value in each equation; just trial and error it

OpenStudy (amistre64):

lets try that y=6 3x + 2y = 6 3x + 2(6) = 6 3x + 12 = 6 3x = -6 x = -2 2x + 3y = −6 2x + 3(6) = −6 2x + 18 = -6 2x = -24 x = -12 since -2 is not the same as -12; we know 6 is not a correct choice

OpenStudy (anonymous):